OpenStudy (amenah8):

how do i graph y=-.75x+5? Please help!

OpenStudy (anonymous):

think of the value -.75 as a fraction and that will help

OpenStudy (anonymous):

what would -.75 be in fraction form

OpenStudy (amenah8):

It should be -3/4.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so your new equation would be y=-3/4x+5 so 3/4= rise/run but its a negative value

OpenStudy (amenah8):

Oh, I see, thank you!

OpenStudy (anonymous):

yeah no problem

OpenStudy (amenah8):

By the way, if the equation was sort of switched around, how would I graph it. This is what I mean: 3x-2y=6